# Long-Context Language Models?

### v2

@@ Appendix F Additional Results
GPT-4 128k 32k 99.9 99.9 98.7 98.3 90.9 84.8 95.4 92.9(3rd) 97.9(2nd)
Command-R (35B) 128k 64k 99.5 99.3 99.3 98.9 97.4 89.6 97.3 96.0(1st) 98.6(1st)
Yi (34B) 200k 16k 98.1 96.9 97.4 95.1 93.0 90.2 95.1 93.8(2nd) 96.4(3rd)
Mixtral (8x7B) 32k 16k 99.4 98.3 98.8 94.3 73.8 42.6 84.5 75.9(5th) 93.1(4th)
Mistral (7B) 32k 4k 98.1 96.2 94.3 85.5 51.1 10.7 72.6 58.8(7th) 86.5(7th)
ChatGLM (6B) 128k 4k 97.5 95.9 91.9 83.6 67.6 50.9 81.2 73.5(6th) 89.0(5th)
LWM (7B) 1M <4k 92.5 92.1 87.6 83.7 84.1 83.4 87.2 85.5(4th) 89.0(6th)
Together (7B) 32k <4k 96.2 89.9 82.3 80.2 0.0 0.0 58.1 40.2(8th) 76.0(8th)
LongChat (7B) 32k <4k 93.3 92.2 81.1 67.3 0.0 0.0 55.7 37.6(9th) 73.7(9th)
LongAlpaca (13B) 32k <4k 74.9 72.2 70.8 53.2 0.0 0.0 45.2 30.7(10th) 59.7(10th)
@@
GPT-4 128k 128k 100.0 100.0 100.0 100.0 100.0 99.6 99.9 99.9(1st) 100.0(1st)
Command-R (35B) 128k 128k 99.9 100.0 99.8 99.5 99.3 89.8 98.1 96.8(2nd) 99.3(2nd)
Yi (34B) 200k 64k 99.8 99.2 98.8 94.5 92.5 76.8 93.6 90.3(3rd) 96.9(3rd)
Mixtral (8x7B) 32k 32k 99.9 99.7 98.4 93.6 86.9 64.2 90.5 85.2(4th) 95.7(4th)
Mistral (7B) 32k 16k 98.9 96.0 92.2 85.0 74.5 0.0 74.4 60.9(5th) 87.9(5th)
ChatGLM (6B) 128k <4k 84.0 81.2 78.0 66.0 38.4 13.0 60.1 48.3(6th) 71.9(7th)
LWM (7B) 1M <4k 84.4 80.1 67.2 52.2 45.9 15.2 57.5 46.5(7th) 68.6(8th)
Together (7B) 32k <4k 89.2 88.8 48.3 16.6 0.0 0.0 40.5 22.8(9th) 58.2(9th)
LongChat (7B) 32k 8k 97.6 93.5 83.4 62.4 0.0 0.0 56.2 37.4(8th) 75.0(6th)
LongAlpaca (13B) 32k <4k 8.5 2.1 18.2 17.0 0.0 0.0  7.6  6.5(10th)  8.8(10th)
@@
GPT-4 128k 64k 99.0 98.3 98.0 95.0 90.1 79.7 93.4 90.4(1st) 96.3(1st)
Command-R (35B) 128k 16k 93.7 93.8 89.3 73.7 54.7 42.0 74.5 65.2(4th) 83.8(2nd)
Yi (34B) 200k 16k 91.4 90.9 86.2 75.3 58.5 43.4 74.3 66.0(3rd) 82.6(3rd)
Mixtral (8x7B) 32k 16k 91.7 82.7 85.7 69.9 80.9 52.7 77.3 72.1(2nd) 82.4(4th)
Mistral (7B) 32k 8k 94.3 90.4 77.4 48.5 42.4 33.7 64.4 53.1(5th) 75.8(5th)
ChatGLM (6B) 128k <4k 79.6 64.3 53.5 43.2 39.6 35.5 52.6 45.4(6th) 59.9(6th)
LWM (7B) 1M <4k 61.3 43.6 38.3 32.8 29.1 29.1 39.0 34.0(7th) 44.0(9th)
Together (7B) 32k <4k 82.3 64.5 43.3 34.8 0.0 0.0 37.5 22.9(9th) 52.1(7th)
LongChat (7B) 32k <4k 74.3 50.7 46.7 51.1 0.0 0.0 37.1 24.8(8th) 49.5(8th)
LongAlpaca (13B) 32k <4k 33.0 27.0 26.0 23.2 0.0 0.0 18.2 12.3(10th) 24.1(10th)
@@
GPT-4 128k 128k 79.0 78.0 76.0 68.0 61.6 59.0 70.3 66.5(1st) 74.0(1st)
Command-R (34B) 128k 64k 67.9 65.4 63.9 62.6 58.1 48.4 61.1 58.2(3rd) 63.9(4th)
Yi (34B) 200k 128k 72.7 71.5 68.4 66.2 64.1 59.9 67.1 65.0(2nd) 69.2(2nd)
Mixtral (8x7B) 32k 64k 77.6 73.0 71.2 64.5 51.2 34.1 61.9 55.0(5th) 68.8(3rd)
Mistral (7B) 32k 32k 72.4 70.0 65.7 57.6 34.4 13.3 52.2 42.5(6th) 62.0(5th)
ChatGLM (6B) 128k 16k 59.1 53.5 50.9 43.5 34.6 27.3 44.8 39.5(7th) 50.1(7th)
LWM (7B) 1M 128k 61.2 57.8 56.7 55.4 54.7 52.6 56.4 55.1(4th) 57.7(6th)
Together (7B) 32k 16k 61.1 58.3 54.2 45.6 0.0 0.0 36.5 24.9(8th) 48.2(8th)
LongChat (7B) 32k 8k 54.5 53.6 47.6 34.0 0.0 0.0 31.6 21.0(10th) 42.3(10th)
LongAlpaca (13B) 32k 16k 57.2 53.5 49.7 39.0 0.0 0.0 33.2 22.3(9th) 44.1(9th)

### v3

@@ Appendix F Additional Results
Gemini-1.5 1M >128K 99.8 99.9 99.6 99.7 99.7 99.6 99.7 99.7(1st) 99.7(1st)
Llama3.1 (8B) 128K 64K 99.9 99.9 99.8 99.6 98.7 92.6 98.4 97.5(3rd) 99.4(2nd)
GLM4 (9B) 1M 64K 99.4 99.2 99.5 99.4 97.3 94.4 98.2 97.5(2nd) 98.9(3rd)
Llama3.1 (70B) 128K 64K 100.0 100.0 99.9 99.6 98.5 78.9 96.1 93.5(5th) 98.8(4th)
GPT-4 128K 32K 99.9 99.9 98.7 98.3 90.9 84.8 95.4 92.9(6th) 97.9(5th)
Command-R-plus (104B) 128K 32K 99.9 99.9 99.4 97.9 89.6 65.7 92.1 87.3(8th) 96.9(6th)
GradientAI/Llama3 (70B) 1M 16K 99.0 98.8 98.3 94.5 91.2 84.9 94.4 92.1(7th) 96.8(7th)
Yi (34B) 200K 16K 98.2 96.8 97.3 95.1 93.0 90.2 95.1 93.8(4th) 96.4(8th)
Qwen2 (72B) 128K 32K 100.0 99.9 99.9 99.4 84.5 48.0 88.6 81.3(11th) 95.9(9th)
Phi3-medium (14B) 128K 8K 98.7 98.5 96.6 95.4 91.9 51.3 88.7 82.6(10th) 94.9(10th)
Mixtral-8x22B (39B/141B) 64K 16K 99.3 99.1 97.7 96.7 89.9 23.8 84.4 74.8(12th) 94.1(11th)
LWM (7B) 1M <4K 92.5 92.1 87.6 83.7 84.1 83.4 87.2 85.5(9th) 89.0(12th)
Mistral-v0.2 (7B) 32K 4K 98.1 96.2 94.3 85.5 51.1 10.7 72.6 58.8(13th) 86.5(13th)
DBRX (36B/132B) 32K 8K 99.4 99.0 93.5 73.4 0.5 0.0 61.0 41.6(14th) 80.3(14th)
Together (7B) 32K <4K 96.2 89.9 82.3 80.2 0.0 0.0 58.1 40.2(15th) 76.0(15th)
LongChat (7B) 32K <4K 93.3 92.2 81.1 67.3 0.0 0.0 55.7 37.6(16th) 73.7(16th)
LongAlpaca (13B) 32K <4K 74.9 72.2 70.8 53.2 0.0 0.0 45.2 30.7(17th) 59.7(17th)
@@
GPT-4 128K 128K 100.0 100.0 100.0 100.0 100.0 99.6 99.9 99.9(2nd) 100.0(1st)
Gemini-1.5 1M >128K 100.0 100.0 100.0 100.0 99.6 100.0 99.9 99.9(1st) 100.0(2nd)
Command-R-plus (104B) 128K 128K 100.0 100.0 100.0 100.0 99.9 97.2 99.5 99.2(3rd) 99.8(3rd)
GLM4 (9B) 1M >128K 99.9 99.6 99.8 99.8 99.6 97.7 99.4 99.1(4th) 99.7(4th)
Qwen2 (72B) 128K 64K 100.0 100.0 100.0 100.0 95.2 79.0 95.7 92.9(5th) 98.5(5th)
Llama3.1 (70B) 128K 64K 100.0 100.0 100.0 100.0 99.9 59.2 93.2 88.3(8th) 98.0(6th)
Llama3.1 (8B) 128K 64K 99.9 99.7 99.7 98.8 97.6 70.4 94.4 90.7(6th) 98.0(7th)
GradientAI/Llama3 (70B) 1M 64K 100.0 100.0 100.0 100.0 99.7 56.2 92.6 87.4(9th) 97.9(8th)
Yi (34B) 200K 64K 99.8 99.2 98.8 94.5 92.5 76.8 93.6 90.3(7th) 96.9(9th)
Mixtral-8x22B (39B/141B) 64K 64K 100.0 100.0 99.8 98.6 96.4 0.0 82.5 70.3(10th) 94.7(10th)
Phi3-medium (14B) 128K 16K 99.6 99.2 98.4 82.1 53.6 26.0 76.5 64.1(11th) 88.9(11th)
Mistral-v0.2 (7B) 32K 16K 98.9 96.0 92.2 85.0 74.5 0.0 74.4 60.9(12th) 87.9(12th)
LongChat (7B) 32K 8K 97.6 93.5 83.4 62.4 0.0 0.0 56.2 37.4(14th) 75.0(13th)
DBRX (36B/132B) 32K 8K 100.0 99.0 72.5 45.8 0.0 0.0 52.9 33.3(15th) 72.5(14th)
LWM (7B) 1M <4K 84.4 80.1 67.2 52.2 45.9 15.2 57.5 46.5(13th) 68.6(15th)
Together (7B) 32K <4K 89.2 88.8 48.3 16.6 0.0 0.0 40.5 22.8(16th) 58.2(16th)
LongAlpaca (13B) 32K <4K 8.5 2.1 18.2 17.0 0.0 0.0  7.6 6.5(17th) 8.8(17th)
@@
Gemini-1.5 1M >128K 97.7 97.7 97.6 98.6 97.3 90.9 96.6 95.8(1st) 97.4(1st)
GPT-4 128K 64K 99.0 98.3 98.0 95.0 90.1 79.7 93.4 90.4(2nd) 96.3(2nd)
Qwen2 (72B) 128K 32K 99.3 98.0 93.1 97.4 78.5 70.3 89.4 84.7(3rd) 94.2(3rd)
Mixtral-8x22B (39B/141B) 64K 32K 97.8 96.9 94.8 88.2 83.3 69.7 88.5 84.0(4th) 92.9(4th)
Command-R-plus (104B) 128K 32K 98.2 96.9 95.2 90.3 82.5 59.5 87.1 81.3(5th) 92.8(5th)
Llama3.1 (70B) 128K 32K 99.9 98.3 98.4 97.1 66.3 39.8 83.3 73.8(6th) 92.8(6th)
Phi3-medium (14B) 128K 16K 90.8 95.1 90.3 82.4 62.1 43.8 77.4 69.3(7th) 85.6(7th)
Yi (34B) 200K 16K 91.4 90.9 86.2 75.3 58.5 43.4 74.3 66.0(8th) 82.6(8th)
GLM4 (9B) 1M 8K 93.5 85.2 78.5 68.1 58.3 49.7 72.2 64.8(9th) 79.6(9th)
GradientAI/Llama3 (70B) 1M 8K 96.4 94.7 74.9 57.0 45.1 41.4 68.3 57.7(10th) 78.8(10th)
Llama3.1 (8B) 128K 8K 97.0 90.1 79.2 54.1 43.5 36.2 66.7 55.5(11th) 77.8(11th)
Mistral-v0.2 (7B) 32K 8K 94.3 90.4 77.4 48.5 42.4 33.7 64.4 53.1(12th) 75.8(12th)
DBRX (36B/132B) 32K 8K 94.5 94.7 73.7 48.7 4.1 0.0 52.6 34.3(13th) 70.9(13th)
Together (7B) 32K <4K 82.3 64.5 43.3 34.8 0.0 0.0 37.5 22.9(16th) 52.1(14th)
LongChat (7B) 32K <4K 74.3 50.7 46.7 51.1 0.0 0.0 37.1 24.8(15th) 49.5(15th)
LWM (7B) 1M <4K 61.3 43.6 38.3 32.8 29.1 29.1 39.0 34.0(14th) 44.0(16th)
LongAlpaca (13B) 32K <4K 33.0 27.0 26.0 23.2 0.0 0.0 18.2 12.3(17th) 24.1(17th)
@@
Gemini-1.5 1M >128K 81.9 75.9 77.8 75.9 77.6 74.1 77.2 76.3(1st) 78.0(1st)
GPT-4 128K 128K 79.0 78.0 76.0 68.0 61.6 59.0 70.3 66.5(4th) 74.0(2nd)
Qwen2 (72B) 128K 64K 80.8 76.9 74.1 66.9 54.5 47.2 66.7 61.0(8th) 72.5(3rd)
GradientAI/Llama3 (70B) 1M >128K 75.6 73.9 72.4 69.9 66.0 59.8 69.6 67.1(3rd) 72.1(4th)
Llama3.1 (70B) 128K 64K 77.2 74.8 72.3 70.4 64.2 47.6 67.8 63.4(7th) 72.1(5th)
GLM4 (9B) 1M >128K 74.7 71.3 71.9 68.5 66.3 63.6 69.4 67.6(2nd) 71.1(6th)
Mixtral-8x22B (39B/141B) 64K 64K 76.6 73.5 71.8 66.5 59.7 40.8 64.8 59.4(9th) 70.2(7th)
Yi (34B) 200K >128K 72.7 71.5 68.4 66.2 64.1 59.9 67.1 65.0(5th) 69.2(8th)
Llama3.1 (8B) 128K 128K 74.1 70.1 67.3 65.8 63.7 58.8 66.6 64.3(6th) 68.9(9th)
Command-R-plus (104B) 128K 64K 73.4 72.3 69.4 65.9 57.0 39.2 62.9 57.6(10th) 68.1(10th)
Phi3-medium (14B) 128K 64K 70.9 67.2 66.1 59.3 54.2 38.0 59.3 54.3(12th) 64.3(11th)
Mistral-v0.2 (7B) 32K 32K 72.4 70.0 65.7 57.6 34.4 13.3 52.2 42.5(13th) 62.0(12th)
LWM (7B) 1M >128K 61.2 57.8 56.7 55.4 54.7 52.6 56.4 55.1(11th) 57.7(13th)
DBRX (36B/132B) 32K 16K 76.0 69.4 59.4 45.0 9.6 0.0 43.2 29.6(14th) 56.9(14th)
Together (7B) 32K 16K 61.1 58.3 54.2 45.6 0.0 0.0 36.5 24.9(15th) 48.2(15th)
LongAlpaca (13B) 32K 16K 57.2 53.5 49.7 39.0 0.0 0.0 33.2 22.3(16th) 44.1(16th)
LongChat (7B) 32K 8K 54.5 53.6 47.6 34.0 0.0 0.0 31.6 21.0(17th) 42.3(17th)